The a 5 per unit $40,000 a. If the sales price is set at $69, how many units must Benson sell to break even? estimates that sales will probably be 8,000 units. What sales price per unit will allow the company to break even? the product heavily and has set the sales price at $70. If sales are 9,000 units, how much can the K Prev 8 of 8 iNext
Answer #1
ques 1
Let "Q" be quantity to be sold
At breakeven profit is 0
S.P * Q - VC * Q -FIXED COSTS = 0
$69*Q- (48+5)* Q - (40000+56000)=0
$69Q-53Q=96000
16Q=96000
Q=6000
so breakeven in units is 6000
ques 2
Let "SP" be sales price
At breakeven profit is 0
S.P * Q - VC * Q -FIXED COSTS = 0
SP*8000- (48+5)*8000 - (40000+56000)=0
8000*(SP-53)=96000
SP-53=12
SP= $    65.00
Ques 3
let expenditure be "x"
70*9000- (48+5)*9000 - (40000+56000) - x=0
153000-96000-x=0
x=57000